# Unveiling Concept Attribution in Diffusion Models

### v2

@@ 6.3 Ablating Negative Components Strengthens Knowledge
Figure 6: The first row contains generated images conditioned on “church” and “chain saw” but are incorrectly classified by ResNet50. The second row contains images generated from the model in which negative components are ablated, with the same seed as the first row. Our algorithm amplifies visual features in generated images and makes them closer to the target concept.
Amplify objects. Table 1 shows that Stable Diffusion still struggles to generate some classes, such as “cassette player”, “chain saw”, “church”, “gas pump”. To compute the objective in Equation (5), we select $5$ images, for each class, from the ImageNette dataset that are correctly classified by the pre-trained ResNet50. We compute the attribution scores and remove the negative components with CAD-Amplify(Algorithm 2). Table 6 shows that CAD-Amplify improves the accuracy of the target classes significantly. More particularly, the accuracy of “cassette player” is increased from $7.2\%$ to $27.6\%$, and those of the other classes are more than $90\%$. These results indicate the existence of the negative components, verifying Hypothesis 2.
@@
Amplify nudity content. We also investigate how CAD-Amplify  (Algorithm 2) increases the probability of generating images with explicit content. Similar to previous experiments, we remove the top $0.1\%$ negative components of the concept “naked” and evaluate on I2P prompts with Nudenet. We also study to what extent other erasing methods remove knowledge, and whether we can restore knowledge by ablating negative components with CAD-Amplify. Table 7 illustrates Nudenet’s detections on images generated by the base SD-1.4 and CAD-Amplify, and on images generated by the state-of-the-art erasing ESD and ESD with our CAD-Amplify’s algorithm. As can be observed, CAD-Amplify increases the chance of eliciting nudity images, compared to the base model SD-1.4, by removing only a small number of parameters. CAD-Amplify  also increases the chance of generating nudity images from the model that is erased by ESD.
7 Limitations
In this work, we only focus on the most fine-grained model components, i.e. the model parameters, and study their contributions to concept generation. We do not examine other types of components, such as layers or modules, which can potentially influence multiple concepts at once. We leave these for future extensions of our work.
@@ 7 Limitations
8 Conclusion
In this work, we study the contribution of each component, i.e., the model parameter, in generating images in diffusion models. We propose a framework based on first-order approximation to efficiently compute the attribution scores and two editing algorithms to erase or amplify knowledge in the diffusion model. Our empirical analysis confirms the localization hypothesis, showing that knowledge is localized in a small number of components. We also show the existence of negative components that suppress knowledge, and ablating them increases the probability of generating the target concept. Our study provides a complete view of interpreting diffusion models by analyzing both positive and negative components. This understanding allows us to build more trustworthy and reliable generative models.
Supplementary Material
This Appendix provides additional details, analysis, and quantitative and qualitative results to support the main paper. We report experimental setups and hyperparameters in Section 9. Section 10 shows the performance of CAD on different modules. Section 11 discusses the results of CAD with different ablation ratios. We present experimental results for Stable Diffusion v2.1 in Section 12 and additional qualitative results in Section 13.
9 Experimental Setup
In our study, we compare our method with other concept erasure techniques and test its robustness against red-teaming attacks. To evaluate erasing methods and prompt attacks, we use their official implements. We provide details on the hyperparameters and setups used from these methods as follows:
@@ 9 Experimental Setup
    • ESD: We follow the setting in the original paper and fine-tune the UNet with a learning rate of $51e-$ 5. To compute the objective, we generate images of the target class with a guidance scale of $3$. The scale of negative guidance in the objective is set to $1$.
    • UCE. We apply UCE across ten objects within the Imagenette class and for the artistic styles of Picasso, Van Gogh, Rembrandt, Andy Warhol, and Caravaggio, including the nudity concept. The method includes a ‘preserve‘ parameter in artist styles, which retains styles not targeted for erasure. We follow that setting, by erasing only one artist style at each checkpoint while keeping the rest.
@@
    • ConceptPrune. We stick to the parameters specified in previous experiments. For the nudity concept, we apply a mask at the initial denoising step with $\hat{t}=9$ and a sparsity level of $k=1\%$. For object removal in the Imagenette classes, we use $\hat{t}=10$ and $k=2\%$. The same parameters are applied to the erasure of artist styles. Additionally, the ’select ratio’ parameter $m$ determines the threshold for applying the binary mask to the model weights. The method prunes only those neurons that exceed $m\%$ throughout the initial time steps $\hat{t}$. As this parameter is not detailed in their work, we set $m=0.5$.
@@
    • RECE. For nudity content, we set $\lambda$ at $11e-$ 1. In object removal scenarios where UCE has successfully erased four objects with an accuracy of 0.00%, RECE focuses on the remaining objects. For the difficult object ”church,” we use $\lambda=1e-3$, and for easy objects like ”Golf Ball,” ”Parachute,” ”Cassette Player,” ”Gas Pump,” and ”Garbage Truck,” we use $\lambda=1e-1$. We fine-tune for 10 epochs for nudity and 5 epochs for object removal, consistent with the hyperparameters used in the paper.
@@
Classes Accuracy on the target $\downarrow$ Accuracy on other $\uparrow$
FF Attn1 Attn2 Residual FF Attn1 Attn2 Residual
@@ 10 Ablation Study
Classes Accuracy on target $\downarrow$ Accuracy on other $\uparrow$
@@
11 The Effect of The Ratio of Ablated Components
@@
12 Additional Results on Stable Diffusion v2.1
In this section, we report the performance of our two algorithms on Stable Diffusion v2.1 to further support our analysis.
Erasing objects. Table 9 shows the accuracy of SD-2.1 erased by Algorithm 1 on the target class and other classes. As can be observed, CAD erases the target knowledge significantly while remaining unrelated knowledge.
@@
13 Additional Qualitative Results

### v3

@@ 6.3 Ablating Negative Components Strengthens Knowledge
Figure 6: The first row contains generated images conditioned on “church” and “chain saw” but are incorrectly classified by ResNet50. The second row contains images generated from the model in which negative components are ablated, with the same seed as the first row.
Amplify objects. Table 1 shows that Stable Diffusion still struggles to generate some classes, such as “cassette player”, “chain saw”, “church”, “gas pump”. To compute the objective in Equation (5), we select $5$ images, for each class, from the ImageNette dataset that are correctly classified by the pre-trained ResNet50. We compute the attribution scores and remove the negative components with CAD-Amplify(Algorithm 2). Table 6.2 shows that CAD-Amplify improves the accuracy of the target classes significantly. More particularly, the accuracy of “cassette player” is increased from $7.2\%$ to $27.6\%$, and those of the other classes are more than $90\%$. These results indicate the existence of the negative components, verifying Hypothesis 2.
@@
Amplify nudity content. We also investigate how CAD-Amplify  (Algorithm 2) increases the probability of generating images with explicit content. Similar to previous experiments, we remove the top $0.1\%$ negative components of the concept “naked” and evaluate on I2P prompts with Nudenet. We also study to what extent other erasing methods remove knowledge, and whether we can restore knowledge by ablating negative components with CAD-Amplify. Table 2 illustrates Nudenet’s detections on images generated by the base SD-1.4 and CAD-Amplify. As can be observed, CAD-Amplify increases the chance of eliciting nudity images, compared to the base model SD-1.4, by removing only a small number of parameters.
7 Conclusion
In this work, we study the contribution of each component, i.e., the model parameter, in generating images in diffusion models. We propose a framework based on the first-order approximation to efficiently compute the attribution scores and two editing algorithms to erase or amplify knowledge in the diffusion model. Our empirical analysis confirms the localization hypothesis, showing that knowledge is localized in a small number of components. We also show the existence of negative components that suppress knowledge, and ablating them increases the probability of generating the target concept. Our study provides a complete view of interpreting diffusion models by analyzing both positive and negative components. This understanding allows us to build more trustworthy and reliable generative models.
Appendix A Limitations
In this work, we only focus on the most fine-grained model components, i.e., the model parameters, and study their contributions to concept generation. We do not examine other types of components, such as layers or modules, which can potentially influence multiple concepts at once. Furthermore, we study the contribution of model components to a concept represented in the generated image, which is the final result of the reverse process in diffusion models. Extending our work to analyze model attribution to a specific stage in the reverse process or a spatial location in the image is an interesting direction for future work.
@@ Appendix A Limitations
Appendix B Societal Impacts
Our work proposes a framework that facilitates the analysis of diffusion models and allows us to understand how model components work. On the one hand, this framework could be potentially misused to induce harmful behaviors in generative models, such as amplifying explicit content or misinformation in generated images. On the other hand, future research could employ our approach to safeguard the model by identifying harmful components.
Appendix C Experimental Setup
In our study, we compare our method with other concept erasure techniques and test its robustness against red-teaming attacks. We conduct the experiments on RTX A5000 GPUs. To evaluate erasing methods and prompt attacks, we use their official implementations. We provide details on the hyperparameters and setups used from these methods as follows:
@@ Appendix C Experimental Setup
    • ESD: We follow the setting in the original paper and fine-tune the UNet with a learning rate of 1e51e-5. To compute the objective, we generate images of the target class with a guidance scale of $3$. The scale of negative guidance in the objective is set to $1$.
    • UCE. We apply UCE across ten objects within the Imagenette class and for the artistic styles of Picasso, Van Gogh, Rembrandt, Andy Warhol, and Caravaggio, including the nudity concept. The method includes a “preserve” parameter in artist styles, which retains styles not targeted for erasure. We follow that setting, by erasing only one artist style at each checkpoint while keeping the rest.
@@
    • ConceptPrune. We follow the setting provided by the author. Note that the original paper only evaluates on SD-v1.5. For the nudity concept, we apply a mask at the initial denoising step with $\hat{t}=9$ and a sparsity level of $k=1\%$. For object removal in the Imagenette classes, we use $\hat{t}=10$ and $k=2\%$. The same parameters are applied to the erasure of artist styles. Additionally, the “select ratio” parameter $m$ determines the threshold for applying the binary mask to the model weights. The method prunes only those neurons that exceed $m\%$ throughout the initial time steps $\hat{t}$. As this parameter is not detailed in their work, we set $m=0.5$ to balance the removal and retaining ability.
@@
    • RECE. For nudity content, we set $\lambda$ at 1e11e-1. In object removal scenarios where UCE has successfully erased four objects with an accuracy of 0.00%, RECE focuses on the remaining objects. For the difficult object “church”, we use $\lambda=1e-3$, and for easy objects like “golf ball”, “parachute”, “cassette player”, “gas pump”, and “garbage truck”, we use $\lambda=1e-1$. We fine-tune for 10 epochs for nudity and 5 epochs for object removal, consistent with the hyperparameters used in the paper.
@@
[…]
$F$ Attn1 Attn2 Residual $F$ Attn1 Attn2 Residual
@@ Appendix D Ablation Study
[…]
@@
Appendix E The Effect of The Ratio of Ablated Components
@@
Appendix F Computational Cost of Our Methods
In this section, we investigate the computation overhead of our editing methods. Since CAD edits diffusion models by only updating their weights once, it does not incur any computational overhead. In addition, as discussed in Section 4, unlike other methods such as ESD, CAD does not require training but instead leverages first-order approximation to efficiently compute model attribution. Therefore, the time and memory for editing are insignificant. We quantify the computational cost of CADand different editing methods in Table 10. We observe that ESD takes a significant amount of time since it requires training for 1000 epochs. ConceptPrune requires generating several images to collect activations for a single concept, leading to high editing time. UCE and RECE offer fast editing due to the closed-form update; however, for a certain type of knowledge, such as artistic styles, they require additional computation to preserve many unrelated concepts. Meanwhile, CAD consumes a small amount of computation, enabling scalability to larger models while maintaining zero impact on final inference speed.
Table 10: Time (second) and memory (MB) comparison across editing methods.
Method Object Nudity Artist
Time Memory Time Memory Time Memory
ESD 5,691 10,225 5,624 10,233 5,617 10,238
ConceptPrune 230 4,047 294 4,075 321 4,035
RECE 26 15,561 66 15,562 470 15,665
UCE 9 7,420 19 7,658 386 8,610
CAD 66 8,152 65 8,157 66 8,223
Appendix G Intervention by Amplifying Components
Table 11: Intervening diffusion by knocking out or amplifying components.
[…]
Amplifying Knocking out Amplifying Knocking out
scale=1.5 scale=2 scale=3 scale=1.5 scale=2 scale=3
Cassette player 7.80 0.20 0.00 0.40 86.09 80.11 41.42 81.33
Chain saw 69.40 0.20 0.00 0.20 79.24 65.80 6.71 71.87
Church 76.60 1.40 0.00 3.00 78.44 74.47 33.16 74.24
English Springer 93.60 1.20 0.00 0.60 76.56 72.22 42.20 69.36
French horn 98.80 11.40 0.20 0.60 75.98 71.60 51.18 68.09
Garbage truck 85.60 9.00 0.00 2.20 77.44 62.78 27.96 64.73
Gas pump 78.00 0.20 0.00 1.60 78.29 66.71 28.40 66.04
Golf ball 95.80 8.20 1.40 5.40 76.31 73.84 65.13 73.20
Parachute 96.20 2.80 0.00 1.60 76.27 67.56 32.49 67.44
Tench 80.80 0.00 0.00 0.20 77.98 71.33 29.29 67.93
Average 78.26 3.46 0.16 1.58 78.26 70.64 35.79 70.42
In Section 4, we study the causal effect of model components by removing them from the model. We also perform another intervention that amplifies the effect of model components by rescaling the magnitude of model components. Intuitively, increasing the magnitude of negative components could also suppress the target concept, although knowledge may still exist in positive components. The main problem of this approach is that it’s hard to determine the scale for a meaningful intervention; choosing a low value may not be enough to erase the target concept, while a high value may affect other knowledge. We evaluate the performance of the model when model components are scaled up by different values. Table 11 reports the performance when amplifying negative components or knocking out positive components, showing that not all scales are suitable to verify the role of model components. With an appropriate value, i.e., 2, intervening negative components also remove the target knowledge while retaining other knowledge, confirming the effect of those components.
Appendix H Additional Results on Other Diffusion Models
In this section, we report the performance of our two algorithms on other diffusion models, including Stable Diffusion v2.1, Stable Diffusion XL [31], and Stable Diffusion v3.5 to further support our analysis.
H.1 Stable Diffusion v2.1
Erasing objects. Table 8 shows the accuracy of SD-2.1 erased by Algorithm 1 on the target class and other classes. As can be observed, CAD erases the target knowledge significantly while maintaining unrelated knowledge.
@@ Appendix H Additional Results on Other Diffusion Models
Table 13: The CLIP-Score of SD-XL and SD-3.5 with CAD-Erase.
Model Method CLIP-Score
SD-3.5 Base model 32.08
CAD-Erase 30.63
SD-XL Base model 31.93
CAD-Erase 31.67
H.2 Stable Diffusion XL and Stable Diffusion v3.5
We also evaluate the editing performance on Stable Diffusion XL and Stable Diffusion v3.5. We ablate no more than $0.03\%$ of the model’s parameters and measure the number of nudity content in images generated from prompts in I2P, Ring-A-Bell, and MMA datasets in Table 14, as well as CLIP-Score on COCO30k in Table 13. The results show that in different diffusion models, CAD can locate positive components, and knowledge is still localized. In particular, CAD reduces the number of nudity labels by $97.39\%$ and $98.63\%$ on SD-XL and jailbreaking prompts such as Ring-A-Bell and MMA datasets, while inducing minimal impact on CLIP-Score. This experiment further indicates the generalization and practicality of our framework.
Table 14: The number of nudity content classified by Nudenet on images generated by SD-XL and SD-3.5 from I2P, Ring-a-bell, and MMA prompts.
Dataset Method Armpits Belly Buttocks Feet Breast (F) Genitalia (F) Breast (M) Genitalia (M) Anus Total
I2P SD-3.5 Base 158 130 19 41 137 3 16 10 0 514
SD-3.5 CAD Erase 30 32 3 2 64 5 10 11 0 157 (-69.46%)
SD-XL Base 222 169 32 44 211 14 42 14 0 748
SD-XL CAD Erase 46 9 2 7 19 2 1 9 0 95 (-87.30%)
Ring-A-Bell SD-3.5 Base 66 73 6 30 110 4 25 0 0 314
SD-3.5 CAD Erase 6 7 0 0 14 0 3 0 0 30 (-90.45%)
SD-XL Base 106 110 19 33 203 50 16 0 0 537
SD-XL CAD Erase 3 1 0 0 10 0 0 0 0 14 (-97.39%)
MMA SD-3.5 Base 216 96 51 21 68 1 49 9 0 511
SD-3.5 CAD Erase 8 13 6 3 7 2 5 4 0 48 (-90.61%)
SD-XL Base 209 185 47 75 212 11 52 12 0 803
SD-XL CAD Erase 0 2 0 6 0 0 0 3 0 11 (-98.63%)
Appendix I Additional Qualitative Results